OpenStudy (anonymous):

if something cost 30$ and the price increased to $40 what is the percentage increase?

OpenStudy (shamim):

40-30=?

OpenStudy (anonymous):

no the percentage increase

OpenStudy (shamim):

Ok

OpenStudy (shamim):

40-30=10 Right?

OpenStudy (shamim):

(10/30)*100%=?

OpenStudy (anonymous):

yes but that does not give me the percentage change

OpenStudy (anonymous):

ty

OpenStudy (shamim):

U r most welcome!!

OpenStudy (anonymous):

wait why would i dived by 30 and not 40?

OpenStudy (anonymous):

@shamim

OpenStudy (skullpatrol):

Because you use the "original" price in the formula. :-)

OpenStudy (anonymous):

oh ok that makes since

OpenStudy (skullpatrol):

$$\Huge \dfrac{percent-of-change}{100} = \dfrac{change-in-price}{original-price}$$